OpenStudy (anonymous):

28a/4a^3= A. 7a^2 B. 7/a^2 C. 1/7a^2

OpenStudy (anonymous):

answer is B. quite clear cut, u need an explanation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay. \[28a /(4a ^{3})= (28/4) * ( a / a ^{3}) = 7 * 1/ a ^{2}= 7/a ^{2}\]

OpenStudy (anonymous):

ty